OpenStudy (anonymous):

Solve: 4(x − 1) = 2(x + 2) + 6

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Just Distribute

OpenStudy (anonymous):

\[4x - 4 = 2x + 4 + 6\] \[2x = 14\]

OpenStudy (anonymous):

Nw Solve For x...

OpenStudy (anonymous):

ALERT X IS NOT 7

OpenStudy (anonymous):

@alanli123

OpenStudy (anonymous):

\[x = \frac{ 14 }{ 2 } = 7\]

OpenStudy (anonymous):

lmao @micahwood50

OpenStudy (anonymous):

not even that funny btw you remind me of : hell yea dirty bass ghetto girl you make me crae